# D-69120 Heidelberg, Germany

### v2

@@ 7 Discussion and Caveats
Figure 12: Constant part of two-power-law parametrization $c$ vs. grain size for CA (left) and SC (right), dotted lines show linear regression for each grain model.
@@ 8 Summary and Outlook
In this paper, we present a numerical study of the impact of grain shape on the radiative torque (RAT) efficiency and its dependence on the magnitude of the radiation field as well as the material and porosity of the dust grains. As a source of radiation we consider O-type stars, an active galactic nuclei (AGN) and the interstellar radiation field (ISRF) typical for the Milky Way. The dust grains are modeled by means of ballistic aggregation and migration (BAM) of monomers of various sizes. The aggregation process is biased to control the size and number of connections of the grain aggregates. We calculate the optical dust grain properties in the discrete dipole approximation using the DDSCAT code considering the refractive indices of pure carbonaceous and silicate grains, respectively. Given the optical properties of the dust aggregates and the radiation energy density we calculate the maximal angular velocity $\omega_{\mathrm{RAT}}$ according to RAT alignment theory. We note again that the conclusions presented here are still preliminary in nature, as stable alignment of grain rotation axis and direction of incoming radiation was assumed and no magnetic fields were included. Our major findings are as follows:
@@
  • We report that the ISRF produced average angular velocities $\omega_{\mathrm{RAT}}=10^{6}\,\mathrm{s}^{-1}$ for both carbon and silicate grains where as the O-type stars and AGN sources produced much larger $\omega_{\mathrm{RAT}}$ in a range of $10^{8}\,\mathrm{s}^{-1}$ to $10^{10}\,\mathrm{s}^{-1}$ for carbon grains and $10^{7}\,\mathrm{s}^{-1}$ to $10^{9}\,\mathrm{s}^{-1}$ for silicate grains.
@@
  • Compared to the O-type stars, for the AGN the angular velocity $\omega_{\mathrm{RAT}}$ of grains did not increase in a similar manner, as photon emission drag forces on the grains increase. This is because of higher dust temperatures in the AGN environment.
@@
  • Considering an entire ensemble of BAM grain aggregates, we propose a new parametrization for the RAT efficiency that follows a broken power-law with a constant part for shorter wavelengths followed by two distinct power-laws to cover longer wavelengths. The proposed parametrization describes the average angular velocities $\omega_{\mathrm{RAT}}$ reached by the ensembles of grains more precisely, than the canonical parametrization derived for compact grains.
@@
  • According to our DDSCAT calculations individual grain aggregates may have sign changes in the RAT efficiency over the entire wavelength regime reducing the net torque integrated over wavelength. However, the appearance of such sign changes does impact the average grain spin-up behavior of the entire ensemble only marginally.
@@
  • The stronger stellar source and the AGN produce angular velocities potentially high enough for rotational disruption of the grains. The results for the angular velocities obtained here could be used to study the rotational disruption of ballistic aggregate grains on the level of individual grains in the future.
@@
Our results for a parametrization with two power-laws for grains composed of monomers, potentially improved by expanding to a parametrization for the distribution of $\omega_{\mathrm{RAT}}$ values, can make detailed studies on grain rotation including grain size, porosity, and material easier by replacing computationally expensive DDSCAT simulations of individual grains. Further improvements could be made by dropping the assumption of unobscured source radiation, including isotropic radiation, surface effects, $\mathrm{H}_{2}$-formation and other additional drag effects. Especially in order to apply the results to grain alignment a general orientation of the grains as well as magnetic fields should be added to complement the analysis.
@@ Acknowledgements.
Appendix A Two power-law parametrization fit parameters
For completeness we provide the full set of fitted parameters of the two-power-law RAT parametrization in table A for CA grains and in table A for SC grains.

### v3

@@ 7 Discussion and caveats
Figure 9: Constant part of the two-power-law parameterization $c$ vs. grain size for CA (left) and SC (right). The dotted lines show the linear regression for each grain model.
It might be convenient to generalize the parameters of the parameterization and give them a functional form dependent on monomer density and grain size. Fig. 9 shows the constant part $c$ of the parameterization as a function of grain size for different monomer densities. The linear increase in $c$ with grain size has a different slope for each monomer density, and the slope roughly doubles from BAM2 to BAM1 and triples from BAM2 to BA. It is therefore possible that for each material there is a functional form of $c$ as a linear function, with the slope depending on the monomer density. To determine whether and how the exponents $q_{1}$ and $q_{2}$ depend on the material or monomer density requires a larger sample of monomer densities and refractive indices (materials).
We have shown that aggregated grains have a different functional form of $Q_{\Gamma}(\lambda)$ than compact grains, and they therefore do not reach angular velocities as high as compact grains would in the same circumstances, with differences of up to an order of magnitude for the smallest grains. However, our results and the canonical parameterization agree for the least dense and largest grains (BA, $250\,\mathrm{nm}$). If the trend of increasing $c$ for a larger grain size and lower monomer density (see Fig. 9) holds true, we would expect $\Gamma_{\mathrm{RAT}}$ to further increase for less dense and larger grains. The relation of a lower monomer density that leads to higher torques makes intuitive sense considering that it increases the total surface area of the grain and therefore the cross section for an interaction with radiation, just as increasing the grain size does. While the torque increases, $\omega_{\mathrm{RAT}}$ does not increase further for larger grains, however, and in some cases, it even decreases again for the largest grains. Aggregate grains larger than those simulated here may potentially not reach much higher angular velocities, as the drag also increases with size.
How much $\omega_{\mathrm{RAT}}$ increases for higher radiation energy densities strongly depends on the optical properties and therefore on the composition of the grains because different drag effects dominate in different environments. The significantly increased surface area and more complex geometry of aggregated grains has to be taken into account.
In environments with highly ionized gas and charged grains, such as an AGN, plasma drag in addition to gas drag and photon drag might additionally play a role for grain dynamics (Draine & Lazarian 1998). However, the charging of BA dust adds an additional level of complexity, where individual monomers may carry different charges and even undergo some stochastic fluctuation over time (see e.g. Matthews et al. 2013; Ma et al. 2013; Matthews et al. 2018). It is beyond the scope of this paper to include the plasma drag in the context of BA. However, the difference of the total grain charge of a BA compared to an equivalent sphere should be smaller than one order of magnitude, as shown in Ma et al. (2013) and Matthews et al. (2018). Hence, we estimate that the canonical parameterization of RAT would still overestimate the resulting angular velocity $\omega_{\mathrm{RAT}}$ for ballistic aggregates in the proximity of an AGN. The dust grains were assumed to have no relative velocity to the gas, but grains close to strong radiation fields might experience linear acceleration by the radiation pressure and reach high drift velocities relative to the surrounding gas. We emphasize that the resulting rotation of the BA grain would then also be influenced by mechanical torques (METs, see e.g. Lazarian & Hoang 2007b; Das & Weingartner 2016; Hoang et al. 2018; Reissl et al. 2023), which work in a way similar to RATs, with gas collisions instead of photon collisions. However, it is still subject of debate how METs enhance the grain rotation in the presence of RATs (see also Lazarian & Hoang 2021).
We emphasize once again that our modeling of RATs did not include the dynamical processes associated with an external magnetic field. As noted in LH07 for an analytical RAT model, the appearance of attractors in the $\Theta-\omega$ phase space is tightly connected to the ratio ${q=\max(Q_{\mathrm{e1}})/\max(Q_{\mathrm{e2}})}$ when an external field is present. This finding was later confirmed by Herranen et al. (2019) and Herranen (2020) based on grain shapes that were modeled as Gaussian ellipsoids. We note, however, that a similar parameter $q$ for a pure mechanical alignment was not found (Das & Weingartner 2016; Reissl et al. 2023). Hence, it remains to be seen whether the ratio $q$ is a reliable proxy to predict the long-term stable alignment with the direction of the magnetic field, considering RATs as a spin-up mechanism of ballistic aggregates. Moreover, in the presence of a magnetic field, this is not only a matter of the efficiency of the RATs. Due to internal dissipation processes in paramagnetic and ferromagnetic materials, a fraction of the angular velocity $\omega$ is transferred into internal heat (see e.g. Davis & Greenstein 1951; Dolginov & Mytrophanov 1976; Purcell 1979). Consequently, the $\omega$ reported in our study may be lower, depending on the magnetic field strength and grain orientation. A similar effect of a reduction in $\omega$ for METs was shown by Reissl et al. (2023). However, to which degree a magnetic field would reduce the maximum possible angular velocity $\omega_{\mathrm{RAT}}$ of BAM aggregates remains to be addressed in forthcoming publications.
@@ 8 Summary and outlook
We presented a numerical study of the impact of the grain shape on the radiative torque (RAT) efficiency and its dependence on the magnitude of the radiation field as well as the material and porosity of the dust grains. As a source of radiation, we considered O-type stars, an AGN, and the ISRF typical for the Milky Way. The dust grains were modeled by means of ballistic aggregation and migration (BAM) of monomers of various sizes. The aggregation process was biased to control the size and number of connections of the grain aggregates. We calculated the optical dust grain properties in the discrete dipole approximation using the DDSCAT code, considering the refractive indices of pure carbonaceous and silicate grains. Based on the optical properties of the dust aggregates and the radiation energy density, we calculated the maximum angular velocity $\omega_{\mathrm{RAT}}$ according to the RAT alignment theory. We note again that the conclusions presented here are still preliminary in nature, as a stable alignment of the grain rotation axis and the direction of incoming radiation was assumed and no magnetic fields were included. Our major findings are listed below.
@@
  • We report that the ISRF produced average angular velocities $\omega_{\mathrm{RAT}}=10^{6}\,\mathrm{s}^{-1}$ for both carbon and silicate grains, whereas the O-type stars and AGN sources produced much larger $\omega_{\mathrm{RAT}}$ in a range of $10^{8}\,\mathrm{s}^{-1}$ to $10^{10}\,\mathrm{s}^{-1}$ for carbon grains and $10^{7}\,\mathrm{s}^{-1}$ to $10^{9}\,\mathrm{s}^{-1}$ for silicate grains.
@@
  • Compared to the O-type stars, the angular velocity $\omega_{\mathrm{RAT}}$ of grains for the AGN did not increase in a similar manner because the photon emission drag forces on the grains increase. The reason is that the dust temperatures in the AGN environment are higher.
@@
  • Considering an entire ensemble of BAM grain aggregates, we proposed a new parameterization for the RAT efficiency that follows a broken power law with a constant part for shorter wavelengths followed by two distinct power laws to cover longer wavelengths. The proposed parameterization describes the average angular velocities $\omega_{\mathrm{RAT}}$ reached by the ensembles of grains more precisely than the canonical parameterization derived for compact grains.
@@
  • According to our DDSCAT calculations, individual grain aggregates may have sign changes in the RAT efficiency over the entire wavelength regime. This reduces the net torque integrated over wavelength. However, the appearance of this sign changes affects the average grain spin-up behavior of the entire ensemble only marginally.
@@
  • The stronger stellar source and the AGN produce angular velocities that are potentially high enough for a rotational disruption of the grains. The results for the angular velocities obtained here could be used to study the rotational disruption of ballistic aggregate grains on the level of individual grains in the future.
@@
Our results for a parameterization with two power laws for grains composed of monomers, potentially improved by expansion to a parameterization for the distribution of $\omega_{\mathrm{RAT}}$ values, can facilitate detailed studies of grain rotation including grain size, porosity, and material by replacing computationally expensive DDSCAT simulations of individual grains. Further improvements could be made by dropping the assumption of unobscured source radiation, including isotropic radiation, surface effects, $\mathrm{H}_{2}$ formation, and other additional drag effects. Especially in order to apply the results to grain alignment, a general orientation of the grains as well as magnetic fields should be added to complement the analysis.
@@ Acknowledgements.
Appendix A Figures
In this section we present the full scope of our simulation results in Figs.A-11.
[…]
figureRATs $\Gamma_{\mathrm{RAT}}$ (solid line) per grain size $a_{\mathrm{eff}}$ calculated with the two-power-law parametrization is shown in comparison to the RATs predicted by the canonical parametrization suggested in LH07 (dotted line) as well as the average of the DDSCAT simulation data (diamonds). Each color represents one of the different sources i.e. ISRF (gray), S1 (dark blue), S2 (light blue), or AGN (red), respectively, utilizing the CA grain material (left column) and the SC material (right column) with monomer-densities BA (top row), BAM (middle row), or BAM2 (bottom row). This figure is to be compared with Fig. 8.
Figure 10: Same as Fig. A but for angular velocity $\omega_{\mathrm{RAT}}$. We emphasize that the canonical parametrization may over-predict the magnitude of $\omega_{\mathrm{RAT}}$ up to one order of magnitude higher for BAM grain aggregates as the two-power-law parametrization.
Figure 11: Same as Fig. 10 but showing the canonical parametrization of LH07 (dotted lines) in comparison with the parametrization for porous grains suggested in TK21 (dashed lines).
Appendix B Two power-law parametrization fit parameters
For completeness we provide the full set of fitted parameters of the two-power-law RAT parametrization in table B for CA grains and in table B for SC grains.